40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6690     Entry time: Sat May 26 01:28:58 2012
Author: Koji 
Type: Update 
Category: General 
Subject: Clipping improvement 

The beam at the dark port was largely clipped.
The arm transmissions (TRX/TRY) were ~0.6 no matter how you try to align.

The PZT1/2 X was scanned and TRX/TRY are now both ~0.9.
This means that the beam was mostly clipped in or after the mode mathing optics.

FPMI contrast is (2.45-0.02)/(2.45+0.02) = 98%
Not so bad although it does not tell anything about clipping before the IFO.

In the next opportunity, we should try to "pin" the beam axis on the arm mirrors using A2L.

1. I am not sure exactly where the clipping was. It means that a small clipping might still be there.

2. I am afraid that the beam is still clipped in the faraday. The beam can be scanned using the steering on the PSL table.

 

Attachment 1: Untitled.png  22 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-